# The infinitesimal structure of manifolds with non-continuous Riemannian metrics

### v1

@@ 4.3 An example of a Sobolev Riemannian metric, neither quasi-Riemannian nor infinitesimally Hilbertian
The functions $\theta_{m}$ converge to a function $\theta$ in $W^{1,p}_{\rm loc}(\mathbb{R}^{d-1})$ as $m\to\infty$.
@@
Then, as $(\psi_{j}-2)\in C_{c}^{\infty}((-1,2)^{d-1})$, the Poincaré inequality together with (4.56) yields that
$\norm{(\psi_{j}-2)}^{p}_{W^{1,p}(\mathbb{R}^{d-1})}$ $\leq C\int_{(-1,2)^{d-1}}|\nabla_{euc}\psi_{j}|^{p}\,{\rm d}\mathcal{L}^{d-1}$
$=C|D_{j}|\int_{B_{r_{j}}(0)}\frac{1}{r_{j}^{p}}|\nabla_{euc}\eta|^{p}\Big{(}\frac{x}{r_{j}}\Big{)}\,{\rm d}\mathcal{L}^{d-1}(x)$
$\leq C2^{(j+1)(d-1)}r^{-p}_{j}\mathcal{L}^{d-1}(B_{r_{j}}(0))|\nabla_{euc}\eta|^{p}_{L^{\infty}}$
$\leq C2^{(j+1)(d-1)}r^{d-1-p}_{j}.$ (4.72)
@@
$\norm{\nabla_{euc}\theta_{m}-\nabla_{euc}\theta_{m^{\prime}}}_{L^{p}((-1,2)^{d-1})}^{p}\leq C\sum_{j=m^{\prime}+1}^{m}2^{(j+1)(d-1)}r^{d-1-p}_{j}\leq C\sum_{j=m^{\prime}+1}^{\infty}2^{(j+1)(d-1)}r^{d-1-p}_{j}.$
Now, (4.50) yields that the right hand side goes to 00 as $m^{\prime}\to\infty$, thus, $(\nabla_{euc}\theta_{m})_{m}$ is a null-sequence in $L^{p}(\mathbb{R}^{d})$. As for all $m$ one has that $\theta_{m}-2\in C_{c}^{\infty}((-1,2)^{d-1},\mathbb{R})$, the Poincaré inequality yields that $\theta_{m}$ converges to a function $\theta$ in $W^{1,p}_{\rm loc}(\mathbb{R}^{d-1})$. ∎
@@
$\theta\equiv 1\ \mathrm{on}\ B_{r_{m}/10}(x).$ (4.76)
@@
$\leq C\frac{\sum_{m\in\mathbb{N}}r^{d-1}_{m}\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\}}{r^{d-1}}\to 0\ \mathrm{as}\ r\to 0.$
@@
$g(x^{\prime},x_{d})=\theta(x^{\prime})Id,\ x^{\prime}\in\mathbb{R}^{d-1},x_{d}\in\mathbb{R}.$ (4.77)
Again $(M,g)$ satisfies (ii) from Assumption 3.1 and $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies (2.1).
@@ 4.3.1 Analysis of quasi-Riemannianity
From the Sobolev embedding theorem, we get that for a $d$-dimensional manifold with a Riemannian metric $g\in W^{1,p}_{\rm loc}$ for some $p>d$, that $g$ is automatically continuous. Then the results we recalled in Subsection 2.2, yield that $(M,\mathsf{d}_{g},g)$ satisfies condition (iii) from Definition 3.6 for $C^{1}$-functions. Theorem 1 shows that this is no longer given if $p<d-1$ and therewith works towards solving Problem 2 in [10], Chapter 3: spazi metrici quasi-riemanniani: the question what regularity assumption on $g$ is necessary such that the $\mathsf{d}_{g}$-slope of a Lipschitz function equals the $g$-norm of the gradient almost everywhere (see also (2.2), Theorem (2.18), and Theorem (6.1) in [9]).
@@ 4.3.2 Analysis of infinitesimal Hilbertianity
From here, we will analyse the infinitesimal structure of the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ and prove Theorem 2. Pick a point $y=(y^{\prime},y_{d})\in M$ such that $y^{\prime}\in S$. We will work in a neighbourhood around $y$ and restrict to local coordinates. Pick a vector $u\in\mathbb{R}^{d}$. We want to understand the metric speed at time 00 of the curve $\gamma^{y,u}:[0,\xi]\to\mathbb{R}^{d},t\mapsto y+tu$, $\xi>0$. Define the null set
@@
Fix an $\varepsilon>0$ and find an $r_{\varepsilon}>0$ such that for $r\in(0,r_{\varepsilon})$, the expressions in (4.65) and (4.66) are smaller than or equal to $\varepsilon$. Now for some $0<r<r_{\varepsilon}$ choose $z\in M$ such that $\mathsf{d}_{euc}(z,y)=\frac{r}{12}$, which implies that any $\frac{4}{3}$-shortest curve transversal to $N_{g}$ (i.e. a curve $\gamma\perp N_{g}$ from $y$ to $z$, such that $L_{g}(\gamma)\leq\frac{4}{3}\mathsf{d}_{g}(y,z)$) from $y$ to $z$ lies inside $B_{r}^{euc}(y)$. Pick such a curve $\gamma$ and assume it to be $1$-$\mathsf{d}_{euc}$-Lipschitz (which can be done if $r$ is small enough, as $\mathsf{d}_{g}$ and $\mathsf{d}_{euc}$ are $\sqrt{2}$-equivalent).
@@
For any $x\in\bigcup_{m\in\mathbb{N}}D_{m}$, write $m(x)$ as the natural number such that $x\in D_{m(x)}$. Now the set $\bigcup_{m\leq J}D_{m}$ is finite so we may write it as $\{x_{l}\}_{l=1}^{k}\subset\mathbb{R}^{d-1}$ for some $k\in\mathbb{N}$. Write $m_{l}:=m(x_{l})$. We will inductively define Lipschitz-continuous curves $\gamma^{(l)}:[0,1]\to\mathbb{R}^{d}$ from $y$ to $z$ for $l=0,\ldots,k$ as follows: Define $\gamma^{(0)}:=\gamma$. Now for $l\geq 1$ define
$t_{l}^{\min}:=\min(t\in[0,1]:\gamma^{(l-1)}_{t}\in\overline{B^{euc}_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R}),$
$t_{l}^{\max}:=\max(t\in[0,1]:\gamma^{(l-1)}_{t}\in\overline{B^{euc}_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R}).$
@@
$w_{l}:=\int_{t_{l}^{\min}}^{t_{l}^{\max}}\dot{\gamma}^{(l-1)}_{t}\,{\rm d}t\in\mathbb{R}^{d},$ (4.82)
@@
Moreover, $\gamma^{(l-1)}=\gamma^{(l)}$ on $[t_{l}^{\min},t_{l}^{\max}]^{c}$, thus
@@
Furthermore, we observe that (4.3.2) yields
@@
for all $l=1,\ldots,k$.
Claim 2. For each $l=1,\ldots,k$, $l^{\prime}\leq l$, $(\gamma^{(l)})^{-1}(\overline{B_{r_{m_{l^{\prime}}}/5}(x_{l^{\prime}})}\times\mathbb{R})$ is a closed interval.
@@
We will prove this by induction. For $l=1$ the claim is true as $(\gamma^{(1)})^{-1}(\overline{B_{r_{m_{1}}/5}(x_{1})}\times\mathbb{R})=[t_{1}^{\min},t_{1}^{\max}]$. Now assume the claim holds true for some $l<k$. Now we have that $(\gamma^{(l+1)})^{-1}(\overline{B_{r_{m_{l+1}}/5}(x_{l+1})}\times\mathbb{R})=[t_{l+1}^{\min},t_{l+1}^{\max}]$. For $l^{\prime}\leq l$ denote by $I^{(l)}_{l^{\prime}}:=(\gamma^{(l)})^{-1}(\overline{B_{r_{m_{l^{\prime}}}/5}(x_{l^{\prime}})}\times\mathbb{R})$. By the construction of $\gamma^{(l+1)}$, (4.55), and (4.56), we have that $t_{l+1}^{\min},t_{l+1}^{\max}\notin I^{(l)}_{l^{\prime}}$. Hence, as $I^{(l)}_{l^{\prime}}$ is a closed interval by the induction hypothesis, it holds that $[t_{l+1}^{\min},t_{l+1}^{\max}]\supset I^{(l)}_{l^{\prime}}$, or $[t_{l+1}^{\min},t_{l+1}^{\max}]\cap I^{(l)}_{l^{\prime}}=\emptyset$. In the first case, we get that $I^{(l+1)}_{l^{\prime}}=\emptyset$ and in the second one it holds $I^{(l+1)}_{l^{\prime}}=I^{(l)}_{l^{\prime}}$. Both are closed intervals, which finishes the induction step and proves the claim. ∎
Finally, we note that for all $l=0,\ldots,k$, $\gamma^{(l)}$ is $1$-$\mathsf{d}_{euc}$-Lipschitz. This is true by assumption for $l=0$. For $l>0$ this follows inductively, using that in $\mathbb{R}^{d}$, $w_{l}=\gamma^{(l-1)}_{t^{\max}_{l}}-\gamma^{(l-1)}_{t^{\min}_{l}}$, hence by the induction hypothesis, $|w_{l}|_{euc}\leq t^{\max}_{l}-t^{\min}_{l}$.
@@
Now (4.81) and (4.3.2), together with the fact that $\gamma^{(k)}$ is $1$-$\mathsf{d}_{euc}$-Lipschitz continuous yield that
@@
$|P^{(d-1)}v_{(1)}|_{euc}$ $=\Big{|}\sum_{l=1}^{k}\int_{(\gamma^{(k)})^{-1}(\overline{B_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R})}P^{(d-1)}\dot{\gamma}^{(k)}_{t}\,{\rm d}t\Big{|}_{euc}$
$\leq\sum_{l=1}^{k}\Big{|}\int_{(\gamma^{(k)})^{-1}(\overline{B_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R})}P^{(d-1)}\dot{\gamma}^{(k)}_{t}\,{\rm d}t\Big{|}_{euc}$
@@
$\leq 2r\frac{\sum_{m\in\mathbb{N}}\frac{r_{m}}{5}|\{x\in D_{m}:\overline{B^{\mathbb{R}^{d-1}}_{r_{m}/5}(x)}\cap\overline{B^{\mathbb{R}^{d-1}}_{r}(y^{\prime})}\}|}{r}\leq 2r\varepsilon,$ (4.90)
@@
for vectors $w_{(1)},w_{(2)}$ with $w_{(1)}+w_{(2)}=z-y$ and $w_{(1)}\in\mathrm{span}(e_{d})$. We want to minimise the right hand side of (4.91) under the thereafter mentioned constraints. This will be done similarly as in the previous example (see from (4.14) onwards).
Write again $z-y=u=(u^{\prime},u_{d})\in\mathbb{R}^{d-1}\times\mathbb{R}$. By potentially applying a linear coordinate transform from $SO(d-1)$, we may assume that $u^{\prime}\in\mathrm{span}(e_{1})$. Then we reduce the right-hand-side of (4.91) by orthogonally projecting $w_{(1)}$ and $w_{(2)}$ onto $\mathrm{span}(e_{1},e_{d})$. Now if $u\in\mathrm{span}(e_{1})$, we get that $w_{(1)}=0$ and hence
@@
As before, together with $\mathsf{d}_{g}\geq\boldsymbol{\delta}_{N_{g}}$ this yields
@@
$\sigma(y,u):=\frac{\sum_{m\in\mathbb{N}}\frac{2r_{m}}{5}|\{x\in D_{m}:\overline{B^{\mathbb{R}^{d-1}}_{r_{m}/5}(x)}\cap\overline{B^{\mathbb{R}^{d-1}}_{|u|_{euc}}(y^{\prime})}\}|}{|u|_{euc}}\to 0,\ \mathrm{as}\ |u|_{euc}\to 0,$
@@
We will next prove an upper bound on $\mathsf{d}_{g}(y,z)$ and work with the regularisation. Fix $\epsilon>0$. For any $\varsigma>0$ and $g_{\varsigma}:=\rho_{\varsigma}*g$, it holds $g_{\varsigma}\leq 2Id$. Hence,
@@
By Lemma 4.7, we can find an $m\in\mathbb{N}$ and an $x\in D_{m}$ such that $|P^{(d-1)}(y)-(x,0)|\leq\epsilon|u|_{euc}$. Now for $\varsigma\in(0,r_{m}/20)$, it holds $g_{\varsigma}=1$ on $\{x\}\times\mathbb{R}$. If $|u_{d}|>|u^{\prime}|_{euc}$, write $b=\frac{|u^{\prime}|_{euc}}{|u_{d}|}$ and note that the concatenation of the curves
@@
Now the argument works similarly as in the previous example: We note that for any $y\in(([0,1]^{d-1})^{c}\times\mathbb{R})\cup(\bigcup_{m\in\mathbb{N}}O_{m}\times\mathbb{R})$ we have that $g$ is continuous in a neighbourhood of $y$ and the metric speed of a curve is given by the $g$-norm of its derivative, wherever it exists. By the results from [28], the minimal weak upper gradient of any $C^{1}$-function $f$ is then also almost everywhere given by the $g$-norm of its $g$-gradient $\nabla_{g}f$. Let $\tilde{\phi}\in C_{c}^{\infty}(\mathbb{R}^{d-1})$ such that $\tilde{\phi}((-2,2)^{d-1})=\{1\}$ and $\tilde{\phi}\equiv 0$ on $((-3,3)^{d-1})^{c}$. Define $\phi\in C^{\infty}(M)$ locally via $\phi:\mathbb{R}^{d-1}\times\mathbb{R},(x^{\prime},x_{d})\mapsto\tilde{\phi}(x^{\prime})$. Now consider the functions
@@
$2G_{1}^{2}+2G_{2}^{2}-(G_{3}^{2}+G_{4}^{2})=1\neq 0$ (4.98)
As in the previous example, this is the main fact that will yield non-infinitesimal Hilberianity. We only have to navigate around the problem that $G_{i}\notin L^{2}(\mathbb{R}^{d},\mathrm{vol}_{g})$. In a final step define a function $\eta\in C_{c}^{\infty}(\mathbb{R})$ such that $\eta(x)=\eta(-x)$ for all $x\in\mathbb{R}$ and such that $\eta\equiv 1$ on $(-1,1)$ and $\eta\equiv 0$ on $(-2,2)^{c}$. For $n\geq 1$ define $\eta_{n}\in C_{c}^{\infty}(\mathbb{R})$ via $\eta_{n}\equiv 1$ on $(-n,n)$, $\eta_{n}(x)=\eta(|x|-n+1)$ for $|x|\geq n$ and for $1\leq i\leq 4$, define the functions $f_{i,n}:\mathbb{R}^{d}\to\mathbb{R}$ via

### v3

@@ 4.3 An example of a Sobolev Riemannian metric, neither quasi-Riemannian nor infinitesimally Hilbertian
The functions $\theta_{m}$ converge to a function $\theta$ in $W^{1,d-1}_{\rm loc}(\mathbb{R}^{d-1})$ as $m\to\infty$.
@@
Then, as $(\psi_{j}-2)\in C_{c}^{\infty}((-1,2)^{d-1})$, the Poincaré inequality together with (4.62) and 4.63 yields that
$\norm{(\psi_{j}-2)}^{d-1}_{W^{1,d-1}(\mathbb{R}^{d-1})}$ $\leq C\int_{(-1,2)^{d-1}}|\nabla_{euc}\psi_{j}|^{d-1}\,{\rm d}\mathcal{L}^{d-1}$
$=C|D_{j}|\int_{B_{r_{j}}(0)}|\nabla_{euc}\eta_{j}|^{d-1}(x)\,{\rm d}\mathcal{L}^{d-1}(x)$
$\leq C2^{(j+1)(d-1)}2^{-j(d+1)}\leq C2^{-j}.$ (4.79)
@@
$\norm{\nabla_{euc}\theta_{m}-\nabla_{euc}\theta_{m^{\prime}}}_{L^{d-1}((-1,2)^{d-1})}^{d-1}\leq C\sum_{j=m^{\prime}+1}^{m}2^{-j}.$
Note that the right hand side goes to 0 as $m^{\prime}\to\infty$, thus, $(\nabla_{euc}\theta_{m})_{m}$ is a null-sequence in $L^{d-1}(\mathbb{R}^{d-1})$. As for all $m$ one has that $\theta_{m}-2\in C_{c}^{\infty}((-1,2)^{d-1},\mathbb{R})$, the Poincaré inequality yields that $\theta_{m}$ converges to a function $\theta$ in $W^{1,d-1}_{\rm loc}(\mathbb{R}^{d-1})$. ∎
@@
$\theta\equiv 1\ \mathrm{on}\ B_{\rho_{m}/10}(x).$ (4.83)
@@
$\leq C\frac{\sum_{m\in\mathbb{N}}r^{d-1}_{m}\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\neq\emptyset\}}{r^{d-1}}\to 0\ \mathrm{as}\ r\to 0.$
@@
$g(x^{\prime},x_{d})=\theta(x^{\prime})\mathrm{Id},\ x^{\prime}\in\mathbb{R}^{d-1},x_{d}\in\mathbb{R}.$ (4.84)
Again $(M,g)$ satisfies Assumption 3.1 and $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ satisfies (2.1).
@@ 4.3.1 Analysis of quasi-Riemannianity
From the Sobolev embedding theorem, we get that for a $d$-dimensional manifold with a Riemannian metric $g\in W^{1,p}_{\rm loc}$ for some $p>d$, that $g$ is automatically continuous. Then the results we recalled in Subsection 2.2, yield that $(M,\mathsf{d}_{g},g)$ satisfies condition (iii) from Definition 3.6 for $C^{1}$-functions. Theorem 1 shows that this is no longer given if $d\geq 3$ and $p\leq d-1$ and therewith works towards solving Problem 2 in [11], Chapter 3: spazi metrici quasi-riemanniani: the question what regularity assumption on $g$ is necessary such that the $\mathsf{d}_{g}$-slope of a Lipschitz function equals the $g$-norm of the gradient almost everywhere (see also (2.2), Theorem (2.18), and Theorem (6.1) in [9]).
@@ 4.3.2 Analysis of infinitesimal Hilbertianity
From here, we will analyse the infinitesimal structure of the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ and prove Theorem 2. Pick a point $y=(y^{\prime},y_{d})\in M$ such that $y^{\prime}\in S$. We will work in a neighbourhood around $y$ and restrict to local coordinates. Pick a vector $u\in\mathbb{R}^{d}$. We want to understand the metric speed at time 0 of the curve $\gamma^{y,u}:[0,\xi]\to\mathbb{R}^{d},t\mapsto y+tu$, $\xi>0$. Define the null set
@@
Fix an $\varepsilon>0$ and find an $r_{\varepsilon}>0$ such that for $r\in(0,r_{\varepsilon})$, the expressions in (4.72) and (4.73) are smaller than or equal to $\varepsilon$. Now for some $0<r<r_{\varepsilon}$ choose $z\in M$ such that $\mathsf{d}_{euc}(z,y)=\frac{r}{12}$, which using that $\mathsf{d}_{g}$ and $\mathsf{d}_{euc}$ are $\sqrt{2}$-equivalent implies that any $\frac{4}{3}$-shortest curve transversal to $N_{g}$ (i.e. a curve $\gamma\perp N_{g}$ from $y$ to $z$, such that $L_{g}(\gamma)\leq\frac{4}{3}\mathsf{d}_{g}(y,z)$) from $y$ to $z$ lies inside $B_{r}^{euc}(y)$. Pick such a curve $\gamma$ and assume it to be 11-$\mathsf{d}_{euc}$-Lipschitz (which can be done if $r$ is small enough, as $\mathsf{d}_{g}$ and $\mathsf{d}_{euc}$ are $\sqrt{2}$-equivalent).
@@
This is in the spirit of the step 2 in Subsection 4.2, with the sole difference that we do not explicitly know $\gamma^{-1}(\{g\equiv\mathrm{Id}\})$, but we can use that $\gamma^{-1}(\{g\equiv\mathrm{Id}\})\subset\gamma^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})\times\mathbb{R})$. Very roughly speaking, we will show that an efficient curve $\gamma$ mainly moves in the $d$-th coordinate direction in $\gamma^{-1}((\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}})\times\mathbb{R})$.
@@
As $\gamma$ is 11-$\mathsf{d}_{euc}$-Lipschitz continuous, (4.88) implies that $\gamma$ is very “short” in $\gamma^{-1}(\bigcup_{m>J}\overline{\Omega_{m}}\times\mathbb{R})$, so, again very roughly speaking, it suffices to study $\gamma$ on $\gamma^{-1}(\bigcup_{m\leq J}\overline{\Omega_{m}}\times\mathbb{R})$ to understand $\gamma$ on $\gamma^{-1}(\bigcup_{m\in\mathbb{N}}\overline{\Omega_{m}}\times\mathbb{R})$.
For any $x\in\bigcup_{m\in\mathbb{N}}D_{m}$, write $m(x)$ as the natural number such that $x\in D_{m(x)}$. Now the set $\bigcup_{m\leq J}D_{m}$ is finite so we may write it as $\{x_{l}\}_{l=1}^{k}\subset\mathbb{R}^{d-1}$ for some $k\in\mathbb{N}$. Write $m_{l}:=m(x_{l})$. In the following, we will essentially argue that the most efficient way to get from $\gamma_{0}$ to $\gamma_{1}$ is to pass the set $\overline{B^{\mathbb{R}^{d-1}}_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R}$ at most once for each $l\in\{1,\ldots,k\}$.
We will inductively define Lipschitz-continuous curves $\gamma^{(l)}:[0,1]\to\mathbb{R}^{d}$ from $y$ to $z$ for $l=0,\ldots,k$ as follows: Define $\gamma^{(0)}:=\gamma$. Now for $l\geq 1$ define
$t_{l}^{\min}:=\min(t\in[0,1]:\gamma^{(l-1)}_{t}\in\overline{B^{\mathbb{R}^{d-1}}_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R}),$
$t_{l}^{\max}:=\max(t\in[0,1]:\gamma^{(l-1)}_{t}\in\overline{B^{\mathbb{R}^{d-1}}_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R}).$
@@
$w_{l}:=\int_{t_{l}^{\min}}^{t_{l}^{\max}}\dot{\gamma}^{(l-1)}_{t}\,{\rm d}t=\gamma_{t_{l}^{\max}}-\gamma_{t_{l}^{\min}}\in\mathbb{R}^{d},$ (4.89)
@@
Moreover, $\gamma^{(l-1)}=\gamma^{(l)}$ on $t_{l}^{\min},t_{l}^{\max}]^{c}$ and, as $\overline{B^{\mathbb{R}^{d-1}}_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R}\subset\mathbb{R}^{d}$ is convex, it holds that $\gamma^{(l)}([t_{l}^{\min},t_{l}^{\max}])\subset\overline{B^{\mathbb{R}^{d-1}}_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R}$. Thus, we have that
@@
Furthermore, we observe that the observations after (4.3.2) together with $\gamma^{(l)}([t_{j}^{\min},t_{l}^{\max}])\subset(\bigcup_{m\leq J}\overline{\Omega_{m}})\times\mathbb{R}\subset(\bigcup_{m>J}\overline{\Omega_{m}})^{c}\times\mathbb{R}$ imply that
@@
for all $l=1,\ldots,k$. For $l^{\prime}\leq l$ denote by $I^{(l)}_{l^{\prime}}:=(\gamma^{(l)})^{-1}(\overline{B^{\mathbb{R}^{d-1}}_{r_{m_{l^{\prime}}}/5}(x_{l^{\prime}})}\times\mathbb{R})$.
Claim 2. For each $1,\ldots,k, ll$$^{\prime}\leq l$, we have that $I^{(l)}_{l^{\prime}}$ is a closed interval.
@@
We will prove this by induction. For $l=1$ the claim is true as $(\gamma^{(1)})^{-1}(\overline{B^{\mathbb{R}^{d-1}}_{r_{m_{1}}/5}(x_{1})}\times\mathbb{R})=[t_{1}^{\min},t_{1}^{\max}]$. Now assume the claim holds true for some $l<k$. Now we have that $(\gamma^{(l+1)})^{-1}(\overline{B^{\mathbb{R}^{d-1}}_{r_{m_{l+1}}/5}(x_{l+1})}\times\mathbb{R})=[t_{l+1}^{\min},t_{l+1}^{\max}]$. By the construction of $\gamma^{(l+1)}$, (4.61), and (4.62), we have that $t_{l+1}^{\min},t_{l+1}^{\max}\notin I^{(l)}_{l^{\prime}}$. Hence, as $I^{(l)}_{l^{\prime}}$ is a closed interval by the induction hypothesis, it holds that $t_{l+1}^{\min},t_{l+1}^{\max}]\supset I^{(l)}_{l^{\prime}}$, or $t_{l+1}^{\min},t_{l+1}^{\max}]\cap I^{(l)}_{l^{\prime}}$$\emptyset$. In the first case, we get that $I^{(l+1)}_{l^{\prime}}$$\emptyset$ and in the second one it holds $I^{(l+1)}_{l^{\prime}}=I^{(l)}_{l^{\prime}}$. Both are closed intervals, which finishes the induction step and proves the claim. ∎
Finally, we note that for all $l=0,\ldots,k$, $\gamma^{(l)}$ is 11-$\mathsf{d}_{euc}$-Lipschitz. This is true by assumption for $l=0$. For $l>0$ this follows inductively, using that in $\mathbb{R}^{d}$, $w_{l}=\gamma^{(l-1)}_{t^{\max}_{l}}-\gamma^{(l-1)}_{t^{\min}_{l}}$, hence by the induction hypothesis, $|w_{l}|_{euc}\leq t^{\max}_{l}-t^{\min}_{l}$.
@@
Now (4.88) and (4.3.2), together with the fact that $\gamma^{(k)}$ is 11-$\mathsf{d}_{euc}$-Lipschitz continuous yield that
@@
$|P^{(d-1)}v_{(1)}|_{euc}$ $=\Big|\sum_{l=1}^{k}\int_{(\gamma^{(k)})^{-1}(\overline{B^{\mathbb{R}^{d-1}}_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R})}P^{(d-1)}\dot{\gamma}^{(k)}_{t}\,{\rm d}t\Big|_{euc}$
$\leq\sum_{l=1}^{k}\Big|\int_{(\gamma^{(k)})^{-1}(\overline{B^{\mathbb{R}^{d-1}}_{r_{m_{l}}/5}(x_{l})}\times\mathbb{R})}P^{(d-1)}\dot{\gamma}^{(k)}_{t}\,{\rm d}t\Big|_{euc}$
[…]
@@
$\leq 2r\frac{\sum_{m\in\mathbb{N}}\frac{r_{m}}{5}|\{x\in D_{m}:\overline{B^{\mathbb{R}^{d-1}}_{r_{m}/5}(x)}\cap\overline{B^{\mathbb{R}^{d-1}}_{r}(y^{\prime})}\neq\emptyset\}|}{r}\leq 2r\varepsilon,$ (4.97)
@@
for vectors $w_{(1)},w_{(2)}$ with $w_{(1)}+w_{(2)}=z-y$ and $w_{(1)}\in\mathrm{span}(e_{d})$. We want to minimise the right hand side of (4.98) under the thereafter mentioned constraints. This will be done similarly as in the previous example (see step 3, from (4.14) onwards).
Write again $z-y=u=(u^{\prime},u_{d})\in\mathbb{R}^{d-1}\times\mathbb{R}$. By potentially applying a linear coordinate transform from SO(d1)SO(d-1), we may assume that $u^{\prime}\in\mathrm{span}(e_{1})$. Then we reduce the right-hand-side of (4.98) by orthogonally projecting $w_{(1)}$ and $w_{(2)}$ onto $\mathrm{span}(e_{1},e_{d})$. Now if $u\in\mathrm{span}(e_{1})$, we get that $w_{(1)}=0$ and hence
@@
As before, together with $\mathsf{d}_{g}=\boldsymbol{\delta}\geq\boldsymbol{\delta}_{N_{g}}$ (see (3.2)) this yields
@@
$\sigma(y,u):=\frac{\sum_{m\in\mathbb{N}}\frac{2r_{m}}{5}|\{x\in D_{m}:\overline{B^{\mathbb{R}^{d-1}}_{r_{m}/5}(x)}\cap\overline{B^{\mathbb{R}^{d-1}}_{|u|_{euc}}(y^{\prime})}\neq\emptyset\}|}{|u|_{euc}}\to 0,\ \mathrm{as}\ |u|_{euc}\to 0,$
@@
We will next prove an upper bound on $\mathsf{d}_{g}(y,z)$ and work with the regularisation. Fix $\epsilon>0$. For any $\varsigma>0$ and $g_{\varsigma}:=\rho_{\varsigma}*g$, it holds $g_{\varsigma}\leq 2\mathrm{Id}$. Hence,
@@
By Lemma 4.8, we can find an $m\in\mathbb{N}$ and an $x\in D_{m}$ such that $|P^{(d-1)}(y)-(x,0)|\leq\epsilon|u|_{euc}$. Now for $\varsigma\in(0,\rho_{m}/20)$, it holds $g_{\varsigma}=1$ on $\{x\}\times\mathbb{R}$. If $|u_{d}|>|u^{\prime}|_{euc}$, write $b=\frac{|u^{\prime}|_{euc}}{|u_{d}|}$ and note that the concatenation of the curves
@@
Now the argument works similarly as in the previous example. We note that for any $y\in(([0,1]^{d-1})^{c}\times\mathbb{R})\cup(\bigcup_{m\in\mathbb{N}}O_{m}\times\mathbb{R})$ we have that $g$ is continuous in a neighbourhood of $y$ and the metric speed of a curve is given by the $g$-norm of its derivative, wherever it exists.
Proposition 4.13.
We will only provide a limited amount of detail, as the idea of the proof is very similar to the one of Proposition 4.6.
Proof.
Let $\tilde{\phi}\in C_{c}^{\infty}(\mathbb{R}^{d-1})$ such that $\tilde{\phi}((-2,2)^{d-1})=\{1\}$ and $\tilde{\phi}\equiv 0$ on $((-3,3)^{d-1})^{c}$. Define $\phi\in C^{\infty}(M)$ locally via $\phi:\mathbb{R}^{d-1}\times\mathbb{R},(x^{\prime},x_{d})\mapsto\tilde{\phi}(x^{\prime})$ Now consider the functions
@@
$2G_{1}^{2}+2G_{2}^{2}-(G_{3}^{2}+G_{4}^{2})=1\neq 0.$ (4.105)
As in the previous example, this is the main fact that will yield non-infinitesimal Hilbertianity. We only have to navigate around the problem that $G_{i}\notin L^{2}(\mathbb{R}^{d},\mathrm{vol}_{g})$. In a final step define a function $\eta\in C_{c}^{\infty}(\mathbb{R})$ such that $\eta(x)=\eta(-x)$ for all $x\in\mathbb{R}$ and such that $\eta\equiv 1$ on (1,1)(-1,1) and $\eta\equiv 0$ on $(-2,2)^{c}$. For $n\geq 1$ define $\eta_{n}$ via $\eta_{n}\equiv 1$ on (n,n)(-n,n), $\eta_{n}(x)=\eta(|x|-n+1)$ for $|x|\geq n$ and for $1\leq i\leq 4$, define the functions $f_{i,n}:\mathbb{R}^{d}\to\mathbb{R}$ via